OpenStudy (anonymous):

The lines graphed are perpendicular. The slope of the red line is -1/3. What is the slope of the green line?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

the line perpendicular to the line with slope m has slope -1/m so slope of green line = -1 / (-1/3) = 3